A car travels at a constant speed of 70 miles/hr. How much time is needed for the car to travel at a distance of 20 miles? Convert the answer to minutes.
Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

70/60=20/x
70x=1200
x=1200/70
x=17.14 minutes

Answer by Alan3354(69443)

A car travels at a constant speed of 70 miles/hr. How much time is needed for the car to travel at a distance of 20 miles? Convert the answer to minutes.
===================
d = r*t
t = d/r
t = 20/70 hours
-----------
1 hour = 60 minutes